shasha12:

What is 21 1/4% expressed as a fraction? 13/200 17/80 21/80 17/40

shasha12:

21/80 ?

shasha12:

@JustSaiyan

dude:

21 and 1/4%?

dude:

Fractions are always over 100

dude:

Maybe that helps you

AP:

21 1/4% is the same as 21.25% Percents are usually "over 100" as @dude suggested Therefore: \[\frac{ 21.25 }{ 100 }\] Move the decimal twice to the right in both the numerator and the denominator: \[\frac{2125}{10000}\] And simplify: \[\frac{ 2125 }{ 10000 } \div 125 = \frac{ 17 }{ 80}\] ~Good Luck (:

shasha12:

thanks